OpenStudy (amonoconnor):

Would someone be willing to check my work on taking this anti-derivative? **I'll post the equations as the 1st comment. Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

The problem: \[f(x) = \frac{p^{6+\sqrt{p}} + 3p^4-2p^2}{p^5}\] \[F(x) = \frac{p^{\sqrt{2}+2}}{\sqrt{2}+2} + 3\ln|p| + p^{-2} + C\]

zepdrix (zepdrix):

f(p), yes? :)

OpenStudy (amonoconnor):

Ah, yes:) Force of habit– my bad!

OpenStudy (amonoconnor):

Other than that, would you say it's good?

zepdrix (zepdrix):

I dunno, gimme a few minutes hehe ^^

zepdrix (zepdrix):

Last two terms look good! Hmmmm, a little confused on that first one... Was the original problem supposed to have a sqrt(2) and not a sqrt(p) in it?

OpenStudy (idku):

\(\large \color{black}{f(p) = \dfrac{p^{6+\sqrt{p}} + 3p^4-2p^2}{p^5}}\) \(\large \color{black}{f(p) = p^1+p^{-4.5}+3p^{-1}-2p^{-3}}\) \(\large \color{black}{F(p) = \dfrac{p^{1\color{red}{+1}}}{{1\color{red}{+1}}}+\dfrac{p^{-4.5\color{red}{+1}}}{-4.5\color{red}{+1}}+3\ln|p|+\frac{2p^{-3\color{red}{+1}}}{-3\color{red}{+1}}}\)

OpenStudy (idku):

Oh, no I am wrong

OpenStudy (idku):

p^(√p) ? that has a closed form integral?

zepdrix (zepdrix):

Yes, apparently it does :) I'm trying to remember the technique for dealing with it hehe

OpenStudy (idku):

zepdrix, is that supposed to be like this?

OpenStudy (idku):

applying the power rule, when x is in the exponent aand the base?? Maybe I don't know something about elementary rules of integration?

zepdrix (zepdrix):

Ya they applied power rule for integration. I'm trying to work it out on paper to understand it. The process probably simplifies down to that rule :o Interesting!

OpenStudy (idku):

It is quite wierd... x^x would yeild some redicoulous results with gamma and other stuff, but when the power is √x we get the power rule..... let's see..

OpenStudy (idku):

\(\color{blue}{\displaystyle \int x^{\sqrt{x}}dx}\)

OpenStudy (amonoconnor):

Yes @zepdrix, it's supposed to be (sqrt.2) in the original function's exponent.

zepdrix (zepdrix):

Oh lol :) Ok that makes things simpler then.

OpenStudy (idku):

√2? nice

zepdrix (zepdrix):

Should it be 1+sqrt(2) leading to 2+sqrt(2)? You have 6+sqrt(2) written right now I guess.

OpenStudy (idku):

\(\large \color{black}{f(p) = \dfrac{p^{6+\sqrt{2}} + 3p^4-2p^2}{p^5}}\) \(\large \color{black}{f(p) = p^1+p^{\sqrt{2}~-5}+3p^{-1}-2p^{-3}}\) \(\large \color{black}{F(p) = \dfrac{p^{1\color{red}{+1}}}{{1\color{red}{+1}}}+\dfrac{p^{\sqrt{2}~-5\color{red}{+1}}}{\sqrt{2}~-5\color{red}{+1}}+3\ln|p|+\frac{2p^{-3\color{red}{+1}}}{-3\color{red}{+1}}}\)

OpenStudy (idku):

then, simplifying is on you

OpenStudy (idku):

but I will still take a look at integral of course...

zepdrix (zepdrix):

Oh yes your work looks good amon :) I was being silly. I forgot that we subtract 5 from the 6.

OpenStudy (amonoconnor):

Good tip for the work-around there too @zepdrix :)

zepdrix (zepdrix):

idku, I think you accidentally subtracted 5 `twice`.

OpenStudy (amonoconnor):

Thank you!

zepdrix (zepdrix):

and maybe did something else silly too hehe

OpenStudy (idku):

it seems correct to me? maybe I am just to tired to actually think...